### v2

# WCCNet: Wavelet-context Cooperative Network for Efficient Multispectral Pedestrian Detection

Abstract
Multispectral pedestrian detection achieves better visibility in challenging conditions and thus is essential to autonomous driving, for which both the accuracy and computational cost are of paramount importance. Most existing approaches treat RGB and infrared modalities equally. They typically adopt two symmetrical backbones for multimodal feature extraction, which ignore the substantial differences between modalities and bring great difficulty for the reduction of the computational cost as well as effective crossmodal fusion. In this work, we propose a novel and efficient framework named Wavelet-context Cooperative Network (WCCNet) that is able to differentially extract complementary features of different spectra with lower computational complexity, and further fuse these diverse features based on their spatially relevant crossmodal semantics. In particular, WCCNet simultaneously explore wavelet context and RGB textures within a cooperative dual-stream backbone, which is composed of adaptive discrete wavelet transform (ADWT) layers and heavyweight neural layers. The ADWT layers extract frequency components for infrared modality, while neural layers handle RGB modality features. Since ADWT layers are lightweight and extract complementary features, this cooperative structure not only significantly reduces the computational complexity, but also facilitates the subsequent crossmodal fusion. To further fuse these infrared and RGB features with significant semantic differences, we elaborately design the crossmodal rearranging fusion module (CMRF), which can mitigate spatial misalignment and merge semantically complementary features in spatially-related local regions to amplify the crossmodal reciprocal information. Results from comprehensive evaluations on KAIST and FLIR benchmarks indicate that WCCNet outperforms state-of-the-art methods with considerable computational efficiency and competitive accuracy simultaneously.
I Introduction
Pedestrian detection[1, 2, 3] is a crucial task widely studied in autonomous driving field owing to its central role in ensuring safety and preventing accidents[4]. Due to the limitations of RGB cameras in capturing pedestrian features under low-light environments, the integration of infrared modality with RGB modality considerably enhance the reliability[5, 6, 7], and thus has attracted growing research interests. Recent years have witnessed rapid development of deep learning-based multispectral pedestrian detection methods.
Detection speed and accuracy are equally critical for time-sensitive autonomous driving scenario, since delayed or low-precision detection will both lead to severe consequences. Therefore, it is vital to explore lightweight and efficient frameworks that also provide competitive accuracy. However, the substantial differences in feature representation and semantic information between modalities pose a great challenge to fast feature extraction and effective crossmodal fusion.
Figure 1: Inference time (ms) vs. $\mathrm{MR}^{-2}$ comparisons with state-of-the-art methods on the KAIST[5] test set, including MSDS[8], ARCNN[9], CIAN[10], TC-DET[11], MBNet[12], MLPD[13], BAANet[14], CMPD[15], and TINet[16]. WCCNet-L, WCCNet-S, and WCCNet-XS are variants of WCCNet with different amounts of parameters. A lower $\mathrm{MR}^{-2}$ represents higher precision. In this way, methods located at lower left corner possess better performance.
Existing multispectral detection methods[17, 9, 18, 8, 13, 12, 19, 20, 16, 11, 14, 15, 10] equally treat the feature extraction of RGB and infrared modalities, and typically apply symmetrical backbones for both RGB and infrared modalities, which will inevitably entail extra computational burden. Therefore, it is quite significant and challenging to find a new network with fewer parameters and high accuracy simultaneously. On the other hand, infrared images and RGB images have clearly distinct features[14, 15]. While RGB images have richer information about textures and colors, infrared images contain valuable geometry details such as shapes, contours, and areas, which will be manifested in high-frequency components. Thus, we explore the integration of wavelets and neural layers to differentially extract features from different modalities. This is inspired by the fact that wavelet-based networks can achieve superior performance in various image applications involving extraction of different frequency features [21, 22, 23].
Furthermore, the obvious visual differences between RGB and infrared modalities also signify a huge semantic distinction between their extracted features, which has posed a challenge to crossmodal fusion. Conventional fusion schemes[18, 8, 9, 13] directly apply simple pixel-wise addition and channel-dimension concatenation, thus they are weak in extracting complementary features. Recent fusion schemes [12, 19, 20, 16] differentiate complementary features from redundant features along channel dimension based on the differences in their numeric values. They assume explicitly or implicitly that features of different modalities are numerically homogeneous. However, it is well-known that the semantic differences between features of different modalities cannot be adequately represented by their numeric differences, especially when there exists a huge semantic gap between them. Since pedestrians in the paired RGB and infrared images are spatially related, we manage to explore semantic correlation between crossmodal features based on their spatial correlation.
In this paper, aiming for faster and more accurate multispectral pedestrian detection, we propose a novel Wavelet-Context Cooperative framework (WCCNet), consists of ADWT-integrated dual-stream backbone and crossmodal rearranging fusion module. The proposed WCCNet differentially extracts features for infrared and RGB modalities with lower computational cost, and effectively aggregates semantically complementary features through the crossmodal context in spatially relevant local regions.
Firstly, we propose a dual-stream backbone composing of CNN layers and adaptive DWT (ADWT) layers, where CNN layers extract spatial-domain features from RGB modality and ADWT layers extract frequency features from infrared modality. The reciprocal relationship between CNN and ADWT layers not only significantly reduces the amount of parameters and inference time, but also improves the extraction of infrared features to complement RGB modality and facilitate the subsequent crossmodal fusion.
Secondly, we propose CrossModal Rearranging Fusion (CMRF) module to fully leverage the semantic interaction of crossmodal features based on their spatial correlation. CMRF first narrows the semantic gap between multimodal features extracted in different domains by the Cross-domain Embedding layer (CE), and spatially align them by the Crossmodal Spatial Alignment layer (CSA) to lessen the spatial misalignments of multimodal features. Based on the mapping relationship between corresponding local regions of the two modalities, CMRF generates on-the-fly attention weights that focus on the locations of pedestrians. These semantic-aware attention weights are utilized to transfer local instance-specific features as complementary information. The proposed CMRF improves the accuracy and robustness of pedestrian detection.
Our main contributions are summarized as follows:
  • A novel framework WCCNet is proposed to achieve considerable detection efficiency and competitive accuracy. The ADWT and neural layers of its dual-stream backbone possess a reciprocal relationship, which conduct differential manners for different modalities to extract complementary features and prepare for further fusion.
  • CMRF is proposed to fuse semantically complementary and spatially relevant features of the two modalities with large discrepancy. Cross-domain Embedding (CE) and Crossmodal Spatial Alignment (CSA) layers are integrated into CMRF module to tackle the cross-domain semantic gap and the crossmodal spatial misalignment problem respectively.
  • We conducted extensive experiments on KAIST and FLIR, two widely-used benchmarks on autonomous driving scenes. Experimental results demonstrate the superiority of WCCNet in detection efficiency, achieving high precision with significant reduction in learnable parameters and computational cost.
II Related Work
Multispectral Pedestrian Detection
Multispectral pedestrian detection has attracted a growing research interest for its essential role in autonomous driving. In the past few years, deep learning-based methods[17, 9, 18, 8, 13, 12, 19, 20, 16, 11, 14, 15, 10] have exhibited state-of-the-art performance and become the mainstream. Most of them typically utilize a dual-stream framework to extract RGB and infrared features both by symmetrical neural layers, while focussing on the multimodal feature extraction and crossmodal feature fusion. Liu et al. [17] employed four fusion architectures which fuse at different stages of feature extraction, and found that the middle fusion produced the highest accuracy. Based on this observation, MSDS[8] further improved the detection accuracy by refining the detection results using cascaded classification subnetwork. Considering position shift problem in weakly aligned multispectral images, AR-CNN[9] designed an aligned region network to adaptively align the multimodal features. Later, for totally unpaired multispectral images, MLPD[13] leveraged multi-label learning to obtain competitive detection accuracy in unpaired conditions. As for crossmodal feature fusion, beyond conventional fusion methods including pixel-wise addition[9] and channel-dimension concatenation[18, 8, 13], MBNet[12] presented a differential modality aware fusion (DMAF) module for better complementary feature extraction. Following the DMAF paradigm of MBNet, there witnessed a series of similar methods that improved upon MBNet and achieved state-of-the-art performance, e.g., PIAFusion[19], RISNet[20], and TINet[16], etc. In addition to feature-level fusion, CMPD[15] conducted decision-level fusion based on Dempster’s combination rule. YOLO-MSLite[7] further explores to improve detection efficiency.
However, these prior works have neglected the intrinsic feature differences between RGB and infrared modalities, and simply utilized symmetrical backbones for RGB and infrared feature extraction. In fact, for infrared modality rich in contour features but deficient in texture features, it is unnecessary to apply heavyweight neural layers at shallow stages, which may results in excessive computational burden and even feature redundancy. Instead, we explore to extract features differentially according to the characteristics of different modalities with lower computational burden.
Meanwhile, crossmodal fusion schemes in prior works [12, 19, 20, 16] assumed features of different spectra are numerically homogeneous. Accordingly, they tended to filter out redundant features with minor numeric differences from multispectral features. However, due to the huge semantic gap between modalities, redundant features of different modalities can present a large discrepancy in their numeric values, which will hinder the crossmodal fusion of these methods. Therefore, in this study, we establish the semantic-aware interaction between crossmodal features based on their spatial consistency, and further exploit their complementary information.
Wavelets in Deep Learning
As an effective tool in time-frequency analysis, wavelets have been widely applied in various signal processing applications [24, 25]. A great deal of researches has been devoted to the combination of fixed wavelets and deep learning networks for both low-level and high-level visual tasks. All these works demonstrate that superior results can be achieved by using wavelet-based deep learning methods. It is natural yet quite non-trivial to design networks learning directly from frequency domain via wavelet transforms on input data [23]. In wavelet scattering networks, the wavelet transform convolutions are cascaded with nonlinear activations and averaging operators [22]. Fujieda et al. [26] proposed a wavelet CNN structure, in which the Haar wavelets introduce multiscale frequency components into network at different stages as a supplement to multiresolution analysis. To suppress the aliasing effect, Li et al. [27] proposed a novel CNN framework named WaveCNet by replacing of the down-sampling operations with fixed DWTs, achieving higher classification accuracy and better noise-robustness. For semantic segmentation of biomedical images, Zhou et al. [28] utilized wavelets to alleviate the learning bias by decomposing images into low-frequency and high-frequency parts.
Recently, there are some researches about fully learnable wavelets[21, 29]. Differing from these literatures, with the goal to prioritize computational efficiency, ADWT layers of WCCNet adopt fixed Haar wavelets modulated by only a few learnable gain coefficients. While fixed Haar wavelets provide computational simplicity, the learnable gain coefficients can strengthen useful frequency components. Thereby WCCNet achieves an optimal balance between accuracy and efficiency. Besides, instead of stacking wavelet with CNN layers in a cascade manner, in which DWT proceeds or follows other types of learning networks [23, 27, 30], WCCNet combines different layers in a parallel and cooperative way. Reciprocal relationship between CNN-extracted RGB features and ADWT-extracted infrared features is further exploited to improve detection accuracy and speed.
Figure 2: Overview framework of WCCNet. The WCCNet consists of three parts: mutli-domain feature extraction backbone, crossmodal rearranging fusion (CMRF) module and detection head. In backbone, WCCNet utilize CNN layers and ADWT layers to extract features of RGB and infrared modalities respectively. CMRF fusion modules are embedded into the first three stages to supplement crossmodal complementary information. The multiscale features of the last three stages are aggregated by feature pyramid mechanism[31], and then are decoded by decoupling detection head[32] to generate final predictions.
Multispectral Fusion Scheme
Fusion between feature maps of different semantic contents is essential for multispectral detection task. Conventional fusion schemes regard input features as a whole and conduct simple numeric combinations, e.g., pixel-wise addition, pixel-wise multiplication, and channel-dimension concatenation, which are incapable of highlighting the crucial parts of input features. In recent years, fusion schemes based on attention mechanisms have emerged. SENet[33] applied channel attention mechanism and boosted its accuracy by weighting each channel based on channel-wise spatial pooling results. Based on SENet, CBAM[34] additionally adopted spatial attention mechanism to spatially weight feature maps with a score map calculated from pixel-wise channel pooling. The variants of the aforementioned channel attention and spatial attention mechanisms were widely used as the fusion schemes of prior multispectral detection works[12, 11, 14, 19, 20, 16].
Recently, there are some works about learning dynamic kernels which may be useful for fusion tasks. CARAFE++[35] and IndexNet[36] were proposed to generate adaptive kernels to aggregate contextual information within related receptive filed. However, for multispectral features with a large semantic gap, these single-modal fusion schemes need pre-concatenation for multimodal input features, and thus can not capture contextual correlations between feature maps of different modalities. Besides, these fusion schemes heavily rely on pixel-wise or channel-wise alignment of input features, thus prone to feature misalignment which are common in practical applications. Instead, the contextual correlations between multimodal features of different modalities and the adaptive spatial feature alignment are explored in this work.
III Proposed Method
The overall architecture of WCCNet is shown in Fig. 2. The WCCNet mainly consists of three parts: (1) wavelet-integrated dual-stream backbone for fast and lightweight multispectral feature extraction; (2) crossmodal rearranging fusion module for semantic-aware crossmodal fusion; (3) decoupling detection head following YOLOX [32] for task-oriented predictions. In the following subsections, we shall mainly introduce the proposed backbone and fusion module.
III-A DWT-integrated Dual-stream Backbone
Overall Structure
For higher computational efficiency and precision, we propose a wavelet-integrated dual-stream backbone to extract features differentially according to the characteristics of different modalities. The proposed backbone consists of two branches: (1) RGB branch employs ordinary learnable convolutions to extract generic features from RGB modality in spatial domain; (2) Infrared branch utilizes wavelet-integrated layers to extract different frequency components for infrared images.
Adaptive DWT layers (ADWT) layer is composed of fixed wavelet kernels and learnable gain factors. The wavelet kernel is implemented by fixed-parameter and separable 1D convolution in two orthogonal directions, and the depth of infrared branch is quite shallower than RGB branch. In this way, the proposed backbone achieves fewer parameters and floating-point operations than ordinary symmetrical CNN backbones.
Let $\mathbf{X}_{R}\in\mathbb{R}^{3\times H\times W}$ and $\mathbf{X}_{T}\in\mathbb{R}^{1\times H\times W}$ denote the input tensors for RGB branch and infrared branch, where $H$ and $W$ represent the height and width of input tensors respectively. The output multiscale features of RGB branch and infrared branch at $j^{th}$ stage are defined as $\mathbf{I}_{R}^{j}\in\mathbb{R}^{C_{R}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ and $\mathbf{I}_{T}^{j}\in\mathbb{R}^{C_{T}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$, where $j\in\{1,2,3\}$, $C_{R}^{j}$ and $C_{T}^{j}$ denote the channel numbers of these two branches at $j^{th}$ stage respectively. Additional two stages are applied to extract higher-level semantic features from the fusion of $\mathbf{X}_{R}$ and $\mathbf{X}_{T}$. Refined multispectral features from the last three stages are adopted for further pedestrian detection, namely $\mathbf{I}_{O}^{i}\in\mathbb{R}^{C^{i}_{O}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}}$, where $i\in\{3,4,5\}$ and $C^{i}_{O}$ denotes the channel numbers at $i^{th}$ stage. Following He et al. [37], residual blocks are employed in RGB branch and the last two stages.
Multiscale ADWT Layer
We now introduce infrared branch stacked by several multiscale ADWT layers in a sequence. Given is the input tensor $\mathbf{X}_{T}^{j-1}\in\mathbb{R}^{C_{TX}^{j-1}\times\frac{H}{2^{j-1}}\times\frac{W}{2^{j-1}}}$ from $(j-1)^{th}$ stage to $j^{th}$ stage. A set of 1D convolutions are performed by ADWT both vertically and horizontally to decompose $\mathbf{X}_{T}^{j-1}$ into a sequence of subdivided frequency components as the infrared feature $\mathbf{I}_{T}^{j}$. The obtained $\mathbf{I}_{T}^{j}$ is of double downsampled size, whose frequency contents vary in channel dimension. Then $\mathbf{I}_{T}^{j}$ will be subsequently fused with $\mathbf{I}_{R}^{j}$ to provide additional frequency-domain details.
The ADWT layer can be regarded as the projection onto an orthogonal basis of small finite waves. Differing from learnable convolutions in CNN, the 1D convolutions conducted by ADWT have fixed weights and consist of two types of kernels, i.e., kernel $\mathbf{h}=(h_{0},h_{1},...,h_{K})$ as high-pass filter and kernel $\mathbf{g}=(g_{0},g_{1},...,g_{K})$ as low-pass filter, where $K$ denotes the kernel length. The corresponding weights for these filters are fixed and can be selected from different wavelet families.
Several issues will arise for us to cope with. Wavelets with long support length are not able to separate close spaced features. Besides, unsymmetrical wavelets could lead to phase distortion in image processing. Therefore, symmetrical wavelets with short support length are chosen in this paper, specifically, Haar wavelet. Moreover, even-sized kernels of symmetrical wavelets will lead to shrinkage in size and potential shift problem. Thus, we symmetrically pad $\mathbf{X}_{T}^{j-1}$ following[38] before convolving $\mathbf{X}_{T}^{j-1}$ with wavelet kernels.
Then these 1D high-pass and low-pass filters convolve $\mathbf{X}_{T}^{j-1}$ by stride $s=2$ in both horizontal and vertical directions, resulting in approximation features $\mathbf{A}^{j}$, horizontal detailed features $\mathbf{DH}^{j}$ and vertical detailed features $\mathbf{DV}^{j}$. We assume that $\mathbf{d}^{j}_{:,m,n}$ and $\mathbf{a}^{j}_{:,m,n}$ are obtained by horizontal convolution with $\mathbf{h}$ and $\mathbf{g}$ on $\mathbf{X}_{T}^{j-1}$ respectively,
$\left\{\begin{aligned} \mathbf{d}^{j}_{:,m,n}&=\sum\nolimits_{k=1}\nolimits^{K}h_{k}\mathcal{P}(\mathbf{X}_{T}^{j-1})_{:,m,{sn+k}}\\ \mathbf{a}^{j}_{:,m,n}&=\sum\nolimits_{k=1}\nolimits^{K}g_{k}\mathcal{P}(\mathbf{X}_{T}^{j-1})_{:,m,{sn+k}}\end{aligned}\right$., (1)
where $m$ and $n$ denote vertical and horizontal coordinates respectively. Note that $m\in\{0,1,...,\frac{H}{2^{j-1}}\}$ and $n\in\{0,1,...,\frac{W}{2^{j}}\}$. Function $\mathcal{P}(\cdot)$ represents symmetric padding in [38]. In the same way, $\mathbf{DH}^{j}$, $\mathbf{DV}^{j}$, and $\mathbf{A}^{j}$ can be obtained by vertical convolution, which are formulated as follows
$\left\{\begin{aligned} \mathbf{DH}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}g_{k}\mathbf{d}^{j}_{:,sm^{\prime}+k,n}\\ \mathbf{DV}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}h_{k}\mathbf{a}^{j}_{:,sm^{\prime}+k,n}\\ \mathbf{A}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}g_{k}\mathbf{a}^{j}_{:,sm^{\prime}+k,n}\end{aligned}\right$., (2)
where $m^{\prime}=(0,1,...,\frac{H}{2^{j}})$ and $n=(0,1,...,\frac{W}{2^{j}})$. Note that ADWT here does not include the highest frequency component $\mathbf{HH}^{j}$, which can be obtained by convolving with kernel $h$ both in horizontal and vertical directions. This choice comes from experimental results in Table IV of Sec. IV-D, which show that the network without $\mathbf{HH}^{j}$ achieves better performance. A possible explanation might be that the highest frequency component has little useful information for pediatrician detection, and its involvement might amplify underlying noise. As for other applications, it might be useful. Finally, we assign learnable gain factors to frequency components $\mathbf{DH}^{j}$, $\mathbf{DV}^{j}$, and $\mathbf{A}^{j}$, so that ADWT has frequency adaptability to datasets. The output tensor $\mathbf{I}_{T}^{j}$ at $j^{th}$ can be represented as follows,
$\mathbf{I}_{T}^{j}=\boldsymbol{\lambda}^{j}[\mathbf{DH}^{j},\mathbf{DV}^{j},\mathbf{A}^{j}],$ (3)
where $[,]$ denotes concatenation at channel dimension. $\boldsymbol{\lambda}^{j}\in\mathbb{R}^{C_{T}^{j}}$ represents gain factors at $j^{th}$ stage adaptive to datasets, which is learned by depthwise $1\times 1$ convolution.
[…]
Analysis on Memory and Calculational Cost
To further analyze the savings in computational and memory cost of the proposed ADWT-integrated subnetwork, we consider a customized task of general settings as follows. Given input feature map $\mathbf{X}_{in}\in\mathbb{R}^{C^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ at $j^{th}$ stage with a spatial resolution of $\frac{H}{2^{j}}\times\frac{W}{2^{j}}$ and $C^{j}$ channels, the output feature map $\mathbf{I}_{out}\in\mathbb{R}^{C^{j+1}\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$ is expected. Note that $C^{j}$ in RGB branch and infrared branch is different, while the former is much bigger than latter. Therefore, for fair efficiency comparison on the same task, the CE module introduced in Sec. III-B is also regarded as one part of the ADWT subnetwork in this analysis.
For the $j^{th}$ stage of CNN subnetwork, the first layer is a $K\times K$ CNN layer for down-sampling with a kernel $\theta_{0}\in\mathbb{R}^{C^{j}\times C^{j+1}\times K\times K}$ and a stride = 2. Then it is followed by a residual block. The residual block contains $2N$ CNN layers, in which $N$ CNN layers are of kernel size $K\times K$ and the other $N$ layers are of kernel size $1\times 1$. The $K\times K$ CNN layer in residual block is supposed to have a kernel $\theta_{1}\in\mathbb{R}^{C^{j+1}\times C^{j+1}\times K\times K}$ with a stride of 1 and a padding size of $\lfloor\frac{K-1}{2}\rfloor$. Following the computational complexity provided by He et al. [39] which ignores the computational cost of bias and batch normalization, the computational cost $P_{CNN}$ and the memory cost $M_{CNN}$ in the forward inference of CNN subnetwork at $j^{th}$ stage can be estimated as
$P_{CNN}\approx\frac{HW}{4^{j+1}}K^{2}\left(C^{j+1}\right)^{2}\left(\frac{C^{j}}{C^{j+1}}+\frac{N(K^{2}+1)}{K^{2}}\right)$ (4)
$M_{CNN}\approx K^{2}C^{j+1}C^{j}+N(K^{2}+1)\left(C^{j+1}\right)^{2}.$
Considering substitute CNN subnetwork with ADWT subnetwork and CE module at $j^{th}$ stage, the CE module introduced in Sec. III-B is also involved for fair comparison. The ADWT subnetwork first convolves the input feature map with paired low-pass and high-pass kernels $g\in\mathbb{R}^{1\times K\times 1}$ and $h\in\mathbb{R}^{1\times 1\times K}$ in vertical and horizontal directions. Then ADWT adopt depthwise $1\times 1$ convolution with learnable gain coefficients to obtain the weighted feature map $\mathbf{I}_{T}^{j}\in\mathbb{R}^{3C^{j}\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$. Then $\mathbf{I}_{T}^{j}$ are convolved by the CE module in III-B with $K\times K$ convolutional layer to obtain the output feature map $\mathrm{I}_{out}$. Therefore, the computational cost $P_{DWT}$ and the memory cost $M_{DWT}$ in the forward propagation of ADWT subnetwork and CE module at $j^{th}$ stage are
$P_{ADWT+CE}\approx\frac{3HW}{4^{j+1}}C^{j}(\underbrace{K+1}_{ADWT}+\underbrace{2K^{2}C^{j+1}}_{CE})$ (5)
$M_{ADWT+CE}\approx\underbrace{2K+3C^{j}}_{ADWT}+\underbrace{6K^{2}C^{j}C^{j+1}}_{CE}.$
Moreover, since the weights of DWT layers are fixed and non-trainable, the ADWT subnetwork requires much lower computational cost in backward propagation, which means less training time. Given the gradient of the output feature map $\delta_{j+1}\in\mathbb{R}^{C^{j+1}\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$, the gradient $\delta_{j}$ of the input feature map and the gradients $\gamma_{j}$ of the CNN kernels are expected. Assuming that there is no learnable bias in CNN layers, the computational cost in the backward propagation of CNN subnetwork at $j^{th}$ stage is
$B_{CNN}\approx\frac{HW}{4^{j}}C^{j+1}K^{2}(\frac{1}{4}NC^{j+1}+C^{j}).$ (6)
In the same way, the computational cost in the backward propagation of ADWT subnetwork at $j^{th}$ stage is
$B_{ADWT+CE}\approx\frac{HW}{4^{j}}(\underbrace{3C^{j}}_{ADWT}+\underbrace{\frac{1}{4}K^{2}C^{j}C^{j+1}}_{CE}).$ (7)
Therefore, in both forward and backward propagation, ADWT subnetwork has much lower computational complexity than the CNN subnetwork at the same stage, which also demonstrated by quantitative experiments in Sec. IV-D.
Figure 3: Detailed Structure of CMRF, consisting of three parts: Cross-domain Embedding (CE) layer to narrow semantic gap, Crossmodal Spatial Alignment (CSA) layer to maintain spatial consistency, and Semantic Rearranging Fusion (SRF) layer to amplify crossmodal complementary information.
III-B Multi-domain Crossmodal Rearranging Fusion
Multimodal fusion of different modalities is important for accurate multispectral pedestrian detection. In our framework, RGB features $\mathbf{I}_{R}^{j}$ in spatial domain and infrared features $\mathbf{I}_{T}^{j}$ in frequency domain have their own characteristic semantics in channel dimension. The simple amplification for numerically different features in existing approaches[12, 19, 20, 16] may ignore the potential semantic gap between those heterogeneous features of different spectra, and turns out to be nearly ineffective as demonstrated in Sec. IV-D.
In spite of semantic diversity, there exists inherent spatial consistency cross multimodal feature maps. In light of this, we propose a novel fusion module named Crossmodal Rearranging Fusion (CMRF). As illustrated in Fig. 3, CMRF consists of three parts: Cross-domain Embedding (CE) layer, Crossmodal Spatial Alignment (CSA) layer, and Semantic Rearranging Fusion (SRF) layer.
Cross-domain Embedding
At the $j^{th}$ stage, infrared features $\mathbf{I}_{T}^{j}$ contains frequency-domain components which is complementary but contrasting to RGB features $\mathbf{I}_{R}^{j}$. Therefore, Cross-domain Embedding Layer (CE) are adopted to embed $\mathbf{I}_{T}^{j}$ into the same domain of features in RGB branch. The embedding layer consists of a few embedding units, specifically two units in our implementation, which conduct $3\times 3$ convolution, batch normalization, and activation function in sequence. Taking one embedding unit as an example, for the input tensor $\mathbf{I}_{T}^{j}$, the embedding unit generates the output tensor $\mathbf{F}_{T}^{j}\in\mathbb{R}^{C_{T}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ of the same size,
$\mathbf{F}_{T}^{j}=\mathrm{LReLU}\left(\alpha\cdot\frac{\mathcal{F}_{conv}(\mathbf{I}_{T}^{j})-\mathrm{E}(\mathcal{F}_{conv}(\mathbf{I}_{T}^{j}))}{\sigma(\mathcal{F}_{conv}(\mathbf{I}_{T}^{j}))}+\beta\right),$ (8)
where $\alpha$ and $\beta$ denote learnable affine parameters, and function $\mathcal{F}_{conv}(\cdot)$ denotes the convolution operation with learnable weights. $\mathrm{E}(\cdot)$ and $\sigma(\cdot)$ denote the mean and the standard deviation respectively. Leaky ReLU[40] (LReLU) is adopted as activation. In this way, for further crossmodal fusion, the $j^{th}$-stage CE layer projects $\mathbf{I}_{T}^{j}$ into semantically embedded infrared features $\mathbf{F}_{T}^{j}$.
Crossmodal Spatial Alignment
Since large amounts of multispectral image pairs are weakly aligned as represented in Fig. 4, it is vital to align feature maps of two spectra to assure crossmodal spatial consistency for subsequent crossmodal fusion. Inspired by [41], we propose Crossmodal Spatial Alignment layer to predict pixel-wise offsets of infrared modality relative to RGB modality for alignment. Differing from self-shifting in [41], the offsets in our framework are predicted from features of different modalities simultaneously and utilized for crossmodal shifting.
Firstly, at $j^{th}$ stage, RGB features $\mathbf{I}_{R}^{j}$ and semantically embedded infrared features $\mathbf{F}_{T}^{j}$ are concatenated in channel dimension and then convolved by a $3\times 3$ convolution layer to get the multimodal representation, i.e., $\mathbf{F}_{M}^{j}\in\mathbb{R}^{C_{M}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$. From this representation, another $3\times 3$ convolution layer is applied to predict pixel-wise offsets. The predicted offset map has the same spatial resolution with $\mathbf{F}_{M}^{j}$, in which channel dimension of pixel $(m,n)$ is 2 corresponding to 2D offsets $\Delta\mathbf{P}(m,n)=(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y)$. To smoothly rearrange the pixels of $\mathbf{F}_{T}^{j}$, pixels of $\mathbf{F}_{T}^{j}$ are not directly shifted by offsets and interpolated to get a spatial aligned feature map $\mathbf{F}_{AT}^{j}$. Instead, the pixels of $\mathbf{F}_{AT}^{j}$ are the weighted sums of corresponding shifted receptive field of $\mathbf{F}_{T}^{j}$, which can be written as
$\mathbf{F}_{AT}^{j}(m,n)=\!\sum_{x=-w}^{+w}\sum_{y=-h}^{+h}\mathbf{U}(x,y)\mathbf{F}_{T}^{j}(m+x+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x,n+y+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y),$ (9)
where the output feature map $\mathbf{F}_{AT}^{j}$ is of the same size with $\mathbf{F}_{T}^{j}$. $\mathbf{U}$ denotes learnable weights of $K\times K$ odd size centered at (0,0), $w$ and $h$ equal to $\frac{K-1}{2}$. $m$ and $n$ are the 2D coordinates in feature maps. Since $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y$ are fractional, the pixels of coordinate $(m+x+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x,n+y+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y)$ in $\mathbf{F}_{T}^{j}$ are approximated via bilinear interpolation of neighboring pixels, namely pixels at $(m+\lfloor x+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x\rfloor,n+\lfloor y+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y\rfloor)$, $(m+\lceil x+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x\rceil,n+\lfloor y+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y\rfloor)$, $(m+\lfloor x+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x\rfloor,n+\lceil y+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y\rceil)$, and $(m+\lceil x+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x\rceil,n+\lceil y+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y\rceil)$.
(a) KAIST Training Set
(b) KAIST Test Set
Figure 4: The statistic of Intersection over Union (IoU) of bounding boxes between paired RGB and infrared images in (a) KAIST training set and (b) KAIST test set with sanitized annotations.
Semantic Rearranging Fusion
Despite the features of infrared modality are accommodated to RGB modality so far, we argue that the aim of multimodal fusion for multispectral pedestrian detection is to adaptively amplify complementary features of pedestrians which are semantically similar across different modalities. Instead of numerically complementary feature extraction in prior works, we propose Semantic Rearranging Fusion (SRF) layer which leverages inherent crossmodal spatial consistency and transfers semantically complementary infrared features in relevant local regions to complement the corresponding RGB features.
We first briefly explain the overall process. Recall that generic RGB features $\mathbf{I}_{R}^{j}\in\mathbb{R}^{C_{R}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ contain global structures and instance textures, and aligned infrared features $\mathbf{F}_{AT}^{j}\in\mathbb{R}^{C_{T}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ are derived from CSA layer as the input multimodal tensors for SRF. Following crossmodal spatial consistency, SRF exploits the relationship between channel-dimension semantic contents of the pixel $\mathbf{I}_{R}^{j}(l)$ located at $l=(m,n)$ of $\mathbf{I}_{R}^{j}$ and the corresponding local region centered at $l=(m,n)$ in $\mathbf{F}_{AT}^{j}$. The channel-dimension semantic contents of $\mathbf{I}_{R}^{j}(l)$ are utilized to generate the corresponding local rearranging weight $\mathbf{W}_{l}$. Then $\mathbf{W}_{l}$ is applied to extract instance-specific complementary features from $\mathbf{F}_{AT}^{j}$ through calculating the weighted sum for pixels of $\mathbf{F}_{AT}^{j}$ spatially relevant to $\mathbf{I}_{R}^{j}$, namely the pixels of corresponding local region centered at $l$ in $\mathbf{F}_{AT}^{j}$.
To be more specific, convolution layers are first performed on $\mathbf{I}_{R}^{j}$ to compress the channel number $C_{R}^{j}$ into a smaller one $C_{RK}^{j}$, resulting in $\mathbf{I}_{RK}^{j}\in\mathbb{R}^{C_{RK}^{j}\times\frac{H}{2^{j}}\times\frac{H}{2^{j}}}$. Then given $C_{RK}^{j}=k_{r}^{2}$ where $k_{r}\in\mathbb{N}^{*}$, the pixel $\mathbf{I}_{RK}^{j}(l)\in\mathbb{R}^{k_{r}^{2}\times 1\times 1}$ located at $l=(m,n)$ is spatially reshaped to get the local rearranging weight $\mathbf{W}_{l}\in\mathbb{R}^{k_{r}\times k_{r}}$. The obtained $\mathbf{W}_{l}$ is spatially normalized by softmax function to avoid altering the mean of the corresponding region in $\mathbf{F}_{AT}^{j}$, which sets the sum of elements in $\mathbf{W}_{l}$ to $1$. The generation of $\mathbf{W}_{l}$ mentioned above can be formally written as
$\mathbf{W}_{l}=\mathcal{S}\left(\mathcal{F}_{conv}(\mathbf{I}_{R}^{j}(l))\right),$ (10)
where $\mathcal{F}_{conv}$ represents convolution operations with learnable weights, and $\mathcal{S}(\cdot)$ refers to softmax function. It is worth noting that $\mathbf{W}_{l}$ generated from $\mathbf{I}_{R}^{j}(l)$ will give high-response values if $\mathbf{I}_{R}^{j}(l)$ suspected to contain pedestrian instances, which is illustrated by the visualization of $\mathbf{W}_{l}$ in Sec. IV-D.
At last, instance-specific complementary features $\mathbf{G}_{T}^{j}$ of infrared modality are extracted by convolving the corresponding local region in $\mathbf{F}_{AT}^{j}$ with $\mathbf{W}_{l}$ centered at $(m,n)$, which can be expressed as
$\mathbf{G}_{T}^{j}(l)=\sum_{x=-k}^{k}\sum_{y=-k}^{k}\mathbf{W}_{l}(x,y)\mathbf{F}_{AT}^{j}(m+x,n+y),$ (11)
where $k=\lfloor\frac{k_{r}}{2}\rfloor$. In this way, detailed features $\mathbf{G}_{T}^{j}$ from infrared modality of the same instance with RGB modality are extracted, and then concatenated with $\mathbf{I}_{R}^{j}$ to semantically complement the corresponding RGB features for further pedestrian detection with higher precision.
IV Experiments
In this section, the performance of the proposed WCCNet is evaluated on several important multispectral pedestrian detection benchmarks and compared with other state-of-the-art methods. Comprehensive ablation experiments are also conducted with quantitative and qualitative illustrations.
IV-A Datasets
Our approach is evaluated on two widely-used multispectral pedestrian detection benchmarks including KAIST[5] and FLIR[6], whose data samples are all collected in autonomous driving scenes. Next, we introduce their detailed information.
KAIST
KAIST[5] contains videos of 95,328 RGB-infrared paired frames with 103,128 bounding boxes in total, which are taken from urban road scenes during day and night to cover changes in diverse lighting conditions, heavy occlusions, and multiple pedestrian scales. The videos are split into two halves following [5], where the first half is used for training, and the latter half for testing. Then those two set of videos are sampled every 20 frames into images to obtain image datasets for training and testing respectively. Among images for training, $20$ percent of them are randomly selected for validation in training phase. Following common annotation settings for a fair comparison with recent related works, the paired annotations in [9] are adopted for training and the cleaned annotations[17] are applied for testing. Despite the application of sanitized annotations, there still exist significant spatial misalignments in KAIST as illustrated in Fig. 4, especially for training set, which calls for methods to counter crossmodal misalignments.
FLIR
FLIR[6] contains targets of richer categories than KAIST, including person, car and bicycle. RGB and infrared image pairs in FLIR are taken from moving vehicles at various situations. Following common annotation settings[42], misaligned RGB-infrared image pairs are removed, resulting in well-aligned 4,129 image pairs for training and 1,013 image pairs for testing.
IV-B Implementation Details
TABLE I: Comparisons with the state-of-the-art methods on KAIST dataset. - means not reported in the corresponding paper. >> means the estimated lower bound for the metrics not reported in the corresponding paper. Red, blue, and green colors denote the first, the second, and the third place respectively.
Methods Time Complexity & Space Complexity $\mathrm{MR}^{-2}$ (All) $\mathrm{MR}^{-2}$ (Reasonable)
Params.(M) FLOPs(G) Time(ms) GPU All Day Night All Day night
Fusion RPN+BF[18] >138.4>138.4 >126.20>126.20 800 - - - - 29.83 30.51 27.62
MSDS[8] >71.56>71.56 >480.03>480.03 228 TITAN X 34.20 32.12 38.83 11.34 10.54 12.94
ARCNN[9] >43.88>43.88 >353.79>353.79 120 1080TI 34.95 34.36 36.12 9.34 9.94 8.38
CIAN[10] >162.5>162.5 >142.19>142.19 66.67 1080TI 35.57 36.06 32.38 14.12 14.78 11.13
TC-DET[11] >63.00>63.00 >156.13>156.13 33.17 1080 - - - 27.11 34.81 10.31
MBNet[12] >99.23>99.23 >139.96>139.96 70 1080TI 31.87 32.39 30.95 8.13 8.28 7.86
MLPD[13] >162.5>162.5 >142.19>142.19 12 2080TI 28.49 28.39 28.69 7.58 7.96 6.95
RISNet[20] - - - - - - - 7.89 7.61 7.08
BAANet[14] >73.96>73.96 - 70 1080TI - - - 7.92 8.37 6.98
CMPD[15] >99.23>99.23 >139.96>139.96 110 1080TI 28.98 28.30 30.56 8.16 8.77 7.31
TINet[16] 61.00 - 87.7 TITAN X - - - 9.15 10.25 7.48
WCCNet 17.50 89.98 10.44 3090 19.05 21.17 13.36 7.37 8.55 4.68
18.87 1080TI
33.61 TITAN X
WCCNet-L 48.35 231.81 20.84 3090 18.41 21.06 12.76 7.28 8.36 5.08
WCCNet-S 9.75 53.29 6.39 3090 19.37 21.97 13.34 7.56 8.63 4.98
WCCNet-XS 7.78 43.56 4.29 3090 22.85 25.51 16.19 10.78 12.36 7.32
(a) All-dataset setting for the whole day
(b) All-dataset setting in the daytime
(c) All-dataset setting at night
(d) Reasonable setting for the whole day
(e) Reasonable setting in the daytime
(f) Reasonable setting at night
Figure 5: Log-scale MR-FPPI curves on KAIST with different experimental settings (all-dataset setting and reasonable setting) and different illumination conditions. The lower the curve, the higher the accuracy represented.
Platform and Parameters
The proposed WCCNet has extended the framework of YOLOX[32] in PyTorch. Note that ADWT layers excel at extracting complementary infrared features, and can reciprocally work with other multiscale CNN backbones via CMRF fusion modules. We utilize NVIDIA RTX 3090 for training, and employ a series of GPUs for testing, as shown in Table I. WCCNet is trained from scratch by Stochastic Gradient Descent (SGD) for 80 epochs with an initial learning rate of 0.001, a weight decay of 0.02, and a batch size of 12. Other hyperparameters follow YOLOX if not mentioned. For preprocessing, RGB images and infrared images are all resized to $640\times 640$ before being input to network. Besides, Mosaic[43] and MixUp[44] are adopted as data augmentation methods for the first three quarters of training epochs. For post-processing, the threshold of Non-Maximum Suppression (NMS) is $0.65$, and the threshold for confidence score of positive targets is set to $0.01$.
Supplementary Details for Network
A lightweight convolutional residual backbone named DarkNet-34 is proposed for WCCNet, which follows DarkNet[45] family and has 34 convolutional layers. Our DarkNet-34 has five stages with $\{1,2,4,3,2\}$ residual blocks respectively. The first three stages are applied in RGB branch for feature extraction and the last two stages are shared by two modalities for further multispectral detection. The number of channels $C_{R}^{j}$ at $j^{th}$ stage is adjustable with a scaling parameter $\tau$ and set to $2^{j}\times 32\tau$, while the number of channels for $j^{th}$-stage embedded feature maps $F_{T}^{j}$ in CE layer of infrared branch is set to $2^{j}\times 16\tau$. In our implementation, the $\tau$ of WCCNet is set to $0.5$. Besides, for comprehensive analysis on the impacts of different amounts of parameters, we also present three additional types of WCCNet, namely WCCNet-L with $\tau=1.0$, WCCNet-S with $\tau=0.25$, and WCCNet-XS with $\tau=0.125$. Smaller $\tau$ means lower computational cost but greater underfitting risk, resulting in faster inference speed but lower accuracy. WCCNet with $\tau=0.5$ is recommended in practical applications as a trade-off between detection accuracy and speed.
For the neck of WCCNet, feature pyramid mechanisms are adopted, including Spatial Pyramid Pooling (SPP)[46] and Feature Pyramid Network (FPN)[31]. Specifically, SPP is applied after the last stage of the backbone for further extraction of high-level semantic information. Besides, FPN is applied for top-down information interaction between features of different scales. The rest part of WCCNet, namely detection head for classification and regression of bounding boxes, remains consistent with YOLOX.
Quantitative Metrics
The metric $\mathrm{MR}^{-2}$ is used for evaluating the accuracy on KAIST. It equals to the area of log-average Miss Rate (MR) over False Positive Per Image (FPPI), where FPPI ranges from $10^{-2}$ to $1$. Besides, the log-average MR-FPPI curves are obtained by varying the threshold of detector from 0 to $1$ to generate different MR-FPPI points and plotting these points in logarithmic-scale axes. For consistent comparison with previous works on FLIR, the COCO-style[47] evaluation metrics are adopted, namely Average Precision (AP) for specific class and mean Average Precision (mAP) for all classes. To evaluate the efficiency of multispectral pedestrian detectors, the memory cost are denoted by the numbers of parameters (Param.), while computational complexity are indicated by the number of floating point operations (FLOPs) theoretically and the inference time (Time) practically. Note that the inference time of our model includes post-processing time. Many previous works may not prioritize the efficiency of their models and have not report corresponding metrics in their papers, thus we estimate their lower bound. The FLOPs metric is estimated based on $640\times 640$ input size.
IV-C Comparison with State-of-the-arts
Comparison on KAIST
We evaluate WCCNet and conduct comparisons with other state-of-the-art methods [18, 8, 9, 10, 11, 12, 13, 14, 15, 20, 16] under reasonable and all-dataset settings following [5], as illustrated in Table I and Fig 5. The reasonable setting means that the evaluation is conducted on a subset of testing set which only consists of pedestrians taller than 55 pixels without heavy occlusion, while all-dataset setting evaluates detectors on the whole testing set.
Accuracy Comparison. The experimental results on detection precision are given in terms of $\mathrm{MR}^{-2}$. As can be observed in Table I, all types of our proposed WCCNet outperforms previous methods under all-dataset and reasonable settings. Besides, the MR-FPPI curves under those two settings are shown in Fig. 5(b) and Fig. 5(c) respectively, which also demonstrates the superiority of WCCNet. Especially for all-dataset setting which is more challenging, the WCCNet-L has taken a lead of $35.38\%$ percent over the best recorded method MLPD[13] with 18.41 $\mathrm{MR}^{-2}$ on the all-day subset. This demonstrates the strong capability of WCCNet to handle small and occlusion pedestrians. Moreover, for detection in low-illumination conditions, the precision results on the night subset under reasonable setting show that WCCNet surpasses previous methods by a large margin no less than $32.66\%$. Therefore, it confirms that our ADWT-integrated infrared branch can extract sufficient representations of pedestrians in low-illumination conditions with remarkably fewer parameters.
Efficiency Comparison. For evaluation on model efficiency, the number of parameters, FLOPs, and the inference time are provided in Table. I. Besides, the Speed vs. $\mathrm{MR}^{-2}$ curves under reasonable setting are shown in Fig. 5(a). It is obvious that, compared with previous works, WCCNet achieves remarkable fast speed with fewer parameters and lower computational cost while maintaining superior accuracy, which also supports the computational complexity analysis in Sec. III-A. Especially, WCCNet-S with only 9.75MM parameters and 53.29GG FLOPs can infer at a speed of $156.49$ frames per second (FPS) while maintain competitive accuracy, which again shows the efficiency of WCCNet and its potential in practical applications.
Comparison on FLIR
TABLE II: Evaluation results on FLIR dataset. Red, blue, and green colors denote the first, the second, and the third place respectively. - means not reported by the paper of corresponding methods.
Method Time (ms) mAP50 AP50 per class
Person Car Bicycle
MMTOD-CG[48] 110 61.40 63.31 70.63 50.26
MMTOD-UNIT[48] 61.54 64.47 70.72 49.43
ThermalDet[49] >40.5>40.5 74.60 78.24 85.52 60.04
ODSC[50] >52.6>52.6 69.62 71.01 82.33 55.53
CFR[42] >16.7>16.7 72.39 74.49 84.91 55.77
GAFF[51] 10.85 72.90 - - -
BU-ATT[52] - 73.10 76.10 87.00 56.10
BU-LTT[52] - 73.20 75.60 86.50 57.40
CMPD[15] 110 69.35 69.64 78,11 59.87
YOLO-MSLite[7] 8.9 75.0 81.70 89.10 54.20
WCCNet 10.50 77.3 84.33 89.51 57.99
WCCNet-L 20.92 78.4 86.16 90.73 57.85
WCCNet-S 6.68 74.7 81.71 88.28 54.06
WCCNet-XS 4.36 70.1 79.94 86.65 43.84
In order to illustrate the versatility of WCCNet when handling other categories, further evaluation is conducted on FLIR which contains multiple types of objects. The experimental results are shown in Table II. Precision results are measured using AP50 for each class and mAP50 over all classes with an IoU threshold of 0.5, which remains the same as in previous works. Table II shows that WCCNet-L gets the best result with $78.4\%$ mAP for overall mAP50, outperforming previous works by at least $4.5\%$. The results demonstrate that our WCCNet is able to generalize well to objects of different categories, not only pedestrians.
IV-D Ablation Study & Further Analysis
In this subsection, we perform ablation experiments on KAIST dataset for detailed quantitative and qualitative analysis of the proposed WCCNet detector. The training settings and other unmentioned hyperparameters are the same across all variants of WCCNet for fair comparison.
Analysis of Dual-stream Multi-domain Backbone
TABLE III: Accuracy and efficiency evaluation of different backbone combinations applied in RGB/Infrared branch.
The conditions with block-out input modality are also evaluated, where ✗  means the block-out layers or modalities.
Input Modalities Backbone Types for Branches Time Complexity & Space Complexity $\mathrm{MR}^{-2}$ (All) $\mathrm{MR}^{-2}$ (Reasonable)
RGB Infrared RGB Branch Infrared Branch Fusion Params. (M) FLOPs (G) Time (ms) All Day Night All Day night
CNN CNN CMRF 25.23 130.82 16.91 19.02 21.77 13.26 7.96 9.54 5.09
ADWT ADWT CMRF 10.00 48.92 4.89 26.56 30.66 17.27 11.73 14.12 6.84
ADWT CNN CMRF 17.50 89.98 10.53 21.37 24.56 14.02 9.08 11.44 4.93
CNN ADWT CMRF 17.50 89.98 10.44 19.05 21.17 13.36 7.37 8.55 4.68
CNN 16.71 64.93 5.70 42.90 35.53 60.58 22.62 15.69 36.61
ADWT 9.77 46.01 4.36 36.00 42.11 21.18 17.57 21.54 9.12
RGB in Daytime
(a) Origin Image
(b) After DWT Transform
(c) CNN at 2nd Stage
(d) ADWT at 2nd Stage
(e) CNN at 3rd Stage
(f) ADWT at 3rd Stage
RGB at Night
infrared in Daytime
infrared at Night
Figure 6: Origin images, DWT transform results, and feature maps of different branches in WCCNet are shown in different columns, which are grouped by different times including daytime and night. The RGB modality is shown in the top row, while the ones of infrared modality is shown in the bottom row. DWT transform results are simply obtained by the weighted sum of $\mathbf{DH}$ and $\mathbf{DV}$ for schematic representation.
Different Backbone Types. To confirm the efficiency of the proposed dual-stream multi-domain backbone, which extracts infrared features by ADWT and RGB features by CNN, the evaluation on the performance of different backbone combinations has been carried out and summarized in Table III. Meanwhile, feature maps corresponding to these combinations are quantitatively visualized in Fig. 6. Specifically, the backbone with dual CNN subnetworks in the first row of Table III simply replaces the origin ADWT subnetwork in infrared branch with CNN subnetwork. For the backbone with dual ADWT subnetworks in the second row of Table III, two symmetrical CE layers are applied both for the two ADWT subnetworks in CMRF module. The original backbone with paired CNN-ADWT subnetworks is shown in the gray row.
Table III reveals an apparent 18.818.8% increase in precision when substituting backbone with paired ADWT-CNN subnetworks of WCCNet by backbone with paired CNN-ADWT subnetworks. Moreover, there are a noteworthy 35.635.6% reduction in inference time and a 7.47.4% increase in accuracy when backbone with dual CNN subnetworks is replaced with backbone with paired CNN-ADWT subnetworks. Therefore, ADWT is more efficient to extract infrared features than CNN. It is also evidenced in Fig. 6 that, compared with the infrared features extracted by CNN subnetwork, the ones extracted by ADWT subnetwork exhibit more distinct and clearer responses to targets both in daytime and at night. The main reason lays in the fact that there are rich contour features inherent in high-frequency components of infrared images which are perfectly suited for ADWT subnetwork to extract. However, a significant decline in accuracy occurs when dual ADWT subnetworks are applied not only to infrared images but also RGB images, which is further supported by the top two rows in Fig. 6 that CNN subnetwork extracted more distinguished features for RGB modality than ADWT. The above observation shows again that ADWT subnetwork owns stronger feature representation ability to extract contour features from infrared images, while CNN subnetwork performs better at extracting rich textures from RGB images. Consequently, it demonstrates the superiority of our dual-stream multi-domain backbone, which differentially extract features in different domain.
Absence of Different Spectrum. Furthermore, we consider the scenarios where a spectrum is blocked out. In such cases, the corresponding branch for the blocked-out spectrum is simply removed, while CMRF modules are also removed except that CE layers is kept for ADWT. As expected, it can be observed in the last two row of Table III that the accuracy decreases significantly regardless of which modality is removed, and the accuracy of infrared modality is higher than that of RGB modality. The possible reason might be that infrared images are less affected by varying illumination and complex weather for pedestrian targets with thermal radiation, which highlights the essentiality of effective feature extraction for infrared modality.
Analysis of ADWT for Infrared Modality
We delve into ADWT for infrared feature extraction in this subsection, conducting analysis on the impacts of different frequency components and the adaptive mechanism of ADWT.
To understand the importance of different frequency components for infrared modality, four combinations of frequency components extracted by ADWT are evaluated. These four combinations are $\{\mathbf{A}^{j}\}$, $\{\mathbf{DH}^{j},\mathbf{DV}^{j}\}$, $\{\mathbf{A}^{j},\mathbf{DH}^{j},\mathbf{DV}^{j}\}$, and $\{\mathbf{A}^{j},\mathbf{DH}^{j},\mathbf{DV}^{j},\mathbf{HH}^{j}\}$ at $j^{th}$ stage respectively, which represent different frequency subbands and the third one includes the first two ones. Adaptive mechanism is applied for every combination for consistent comparison. The detection precision of different combinations is shown in the top six rows of Table IV, The accuracy of using $\{\mathbf{A}^{j}\}$ alone and using $\{\mathbf{DH}^{j},\mathbf{DV}^{j}\}$ alone is similar, indicating that $\{\mathbf{A}^{j}\}$ and $\{\mathbf{DH}^{j},\mathbf{DV}^{j}\}$ subbands both are indispensable. Besides, when HH components are reserved by ADWT layer, there is a $2.2\%$ reduction in accuracy since HH components are vulnerable to the influence of noises. Consequently, the combination of $\{\mathbf{A}^{j},\mathbf{DH}^{j},\mathbf{DV}^{j}\}$ is chosen to be extracted by ADWT, which outperforms other combinations in precision.
Then we evaluate the detection performance with/without the adaptive mechanism with learnable gain factors $\boldsymbol{\lambda}$. The adaptive mechanism of ADWT applies different weights to different frequency components, which can be regarded as weight function with learnable parameters $\boldsymbol{\lambda}$. We compare vanilla direct concatenation and adaptive mechanism in Table V. The results show that accuracy decay occurs when adaptive mechanism is replaced by direct concatenation.
TABLE IV: Accuracy comparison on different combinations of different infrared frequency components extracted by ADWT.
Frequency Components $\mathrm{MR}^{-2}$ (All) $\mathrm{MR}^{-2}$ (Reasonable)
$\mathbf{A}$ $\mathbf{DH}$ $\mathbf{DV}$ $\mathbf{HH}$ All Day Night All Day Night
19.30 21.80 13.76 7.76 9.24 5.05
19.66 22.33 14.30 8.02 9.42 5.39
19.05 21.17 13.36 7.37 8.55 4.68
19.18 21.96 13.39 7.53 8.75 4.39
TABLE V: Accuracy comparison on using or not using adaptive mechanisms with learnable gain factor $\boldsymbol{\lambda}^{j}$ for ADWT at $j^{t}h$ stage.
Adaptive Mechanism $\mathrm{MR}^{-2}$ (All)
All Day Night
w/o 19.21 21.98 13.42
w/ 19.05 21.17 13.36
Effectiveness of Wavelet-integrated Backbone
To evaluate the efficiency of wavelet-integrated backbone, we proceed with further experiments on the performance of backbones with ADWT applied in different stages. Specifically, the variants of wavelet-integrated backbone gradually substitutes CNN layers by ADWT layers from stage 1 to stage 5. The precision and computational complexity of these variants are evaluated on multiple classification datasets, namely MNIST[53], FashionMNIST[54], and CIFAR-10[55]. The experimental results are shown in Table VI. As ADWT layers gradually replace CNN layers from shallow layers to deep layers, the numbers of parameters and floating point operations progressively decrease, along with the precision.
Grayscale Images. The classification precision slightly decreases when the backbone varies from the highest to the lowest complexity, $0.89\%$ for MNIST and $5.54\%$ for FashionMNIST. At meanwhile, the reduction in network complexity is substantial, almost $89.9\%$ reduction for FLOPs and $91.2\%$ reduction for parameters. MNIST and FashionMNIST datasets contain grayscale images and thus rich contour features, which will be well extracted by ADWT layers in frequency domain.
RGB Images. As for CIFAR-10 containing RGB images with abundant texture features and colors, the variants of backbone which substitute ADWT layers for CNN layers in deeper network exhibit a $32.08\%$ degradation in precision, but still remain $62.06\%$ precision with a tenfold decrease in computational load. The experimental results on RGB images indicate that the Wavelet-integrated backbone is more suitable to extract grayscale features than RGB features.
TABLE VI: Evaluation on ADWT-integrated backbone variants which interchange CNN layers with ADWT layers at different stages. The accuracy and computational complexity are evaluated on classification benchmarks.
The Stages that CNN Interchanged with ADWT
Stage 1 Stages 1-2 Stages 1-3 Stages 1-4 Stages 1-5
Datasets Precision
MNIST[53] 0.9964 0.9964 0.995 0.9936 0.9875
FashionMNIST[54] 0.9382 0.9368 0.9329 0.9172 0.8862
CIFAR-10[55] 0.9137 0.9142 0.8885 0.8173 0.6206
Metrics Efficiency
FLOPs (G) 1.920 1.600 0.993 0.544 0.194
Params. (M) 22.2 22.0 20.4 16.0 1.96
Early-Cat
(a) Complex Scenario in Daytime
(b) Complex Scenario at Night
Mid-Cat
DMAF
Inter-MA
Late-Cat
CMRF
Figure 7: Comparison on the detection results of WCCNet with different fusion schemes, which are overlaid on RGB/infrared image pairs (in columns). The corresponding attention maps of the detection head are also drawn in the third column. In difficult scenarios, there are occlusions, interfering targets, and small-scale pedestrians. Blue, red and purple boxes stand for true positives, false positives, and false negatives, respectively.
TABLE VII: Accuracy and efficiency evaluation of different fusion methods, along with ablation study on the proposed components of CMRF.
Fusion Scheme
for WCCNet
Efficiency $\mathrm{MR}^{-2}$ (All)
Params. (M) Time (ms) All Day Night
Early Fusion
Early-Cat[18] 16.71 6.91 20.99 23.27 15.04
Middle Fusion
Mid-Cat[13] 16.76 6.14 20.93 23.54 14.04
DMAF[12] 17.51 8.38 20.64 23.29 14.75
Inter-MA[16] 17.13 7.24 21.54 24.80 14.20
Late Fusion
Late-Cat[18] 16.80 6.02 22.67 25.04 17.98
Proposed CMRF
SRF CE CSA Params. (M) Time (ms) All Day Night
17.16 8.63 20.20 22.85 13.84
17.18 8.66 19.53 22.34 12.65
17.49 10.45 19.90 22.54 14.36
17.50 10.44 19.05 21.17 13.36
Analysis of Crossmodal Rearranging Fusion
Quantitative Comparisons with Different Fusion Schemes. Table VII shows ablation studies on comparison between our proposed CMRF and other various fusion schemes. The different fusion schemes are all implemented within WCCNet framework. The Early Concatenation (Early-Cat) and Late Concatenation (Late-Cat) fusion schemes only concatenate multimodality feature maps at the first and the last stage of the backbone along the channel dimension respectively, and get the lowest precision. The Middle Concatenation (Mid-Cat) fusion scheme concatenates RGB and infrared features along the channel dimension at every stage of the backbone, and thus obtain higher accuracy with a minor decline in $MR^{-2}$. These channel-dimension concatenation-based fusion schemes neglect the interaction between multimodal features which is essential for extracting complementary information and filtering out redundant information, hence unsurprisingly they get the lowest precision.
The middle fusion schemes designed for crossmodal feature interaction including DMAF, Inter-MA, and CMRF, are also evaluated. It is worth mentioning that the features for middle fusion are all extracted after CE layer. The DMAF[12] and Inter-MA[16] both obtain complementary information by subtracting feature maps of different modalities along the channel dimension to extract numerically differential features. Instead, the proposed CMRF extracts complementary information through mining semantic-aware features in spatially relevant local regions shared by different modalities.
Daytime
(a) RGB
(b) Infrared
(c) Dynamic weight map
Night
Figure 8: RGB images, infrared images and dynamic weight maps of CMRF are shown in the above three columns respectively. These images are grouped by different times including daytime and night. The highlighted region in red color is the regions of interest for CMRF.
Among the above fusion schemes, CMRF achieves the highest accuracy with competitive inference time, outperforming the second-best DMAF by $7.7\%$ in accuracy.
Qualitative Comparisons with Different Fusion Schemes. We further conduct qualitative evaluation on different fusion schemes. The detection results of WCCNet compared with different fusion schemes are shown in Fig. 7, along with their attention maps of detection head after fusion. It can be seen that, in both day and night, WCCNet has the least amount of falsely detected or missed pedestrians. Even in complex scenarios where pedestrians are occluded or of small scale, WCCNet still locates them accurately.
Analysis of Different Components of CMRF. The influence of different components of CMRF is investigated in Table VII. The baseline is set to that only Semantic Rearranging Fusion(SRF) is applied in WCCNet, while achieving $20.20$ $MR^{-2}$. The accuracy is improved to $19.53$ $MR^{-2}$ when Cross-domain Embedding(CE) layers are added to narrow the semantic gap of multimodal features before being fused by SRF. Lastly, the Cross-domain Spatial Alignment(CSA) is adopted, and then the accuracy further reaches $19.05$ $MR^{-2}$ as features of different spectra are aligned. From this, we can conclude that each component of CMRF indeed contributes to the final accuracy.
Visual Illustration for CMRF. The effectiveness of CMRF is shown in Fig. 8, where the focused regions of context-specific features are visualized. The dynamic weight map in Fig. 8 is implemented by mapping local rearranging weight $\mathbf{W}_{l}$ of all localizations $l$ onto the corresponding images to get heatmaps. As shown in Fig. 8, even when the pedestrians are not clearly visible in any single modality, CMRF effectively focuses on the regions relevant to pedestrians. This observation demonstrates that the complementary features of pedestrians in spatially-related local regions for different modalities are effectively aggregated by CMRF.
V Conclusion
In this paper, we have proposed a novel framework WCCNet to achieve efficient multispectral pedestrian detection for autonomous driving. Based on integrated wavelet with lower computational complexity, WCCNet has boosted the inference speed by learning multi-domain representations for different spectra while maintaining competitive accuracy. A novel fusion scheme CMRF has been proposed to merge the spatially relevant local regions shared by different spectra with adaptive weights learned on-the-fly. Extensive experiments on KAIST and FLIR datasets have been conducted and substantiate the aforementioned superiority of WCCNet over several state-of-the-art methods.